Freshman Seminar 21n: Elliptic Curve Projects

William Stein April 29, 2003

Today you will hand in a more polished draft of your project. Your assignment for the week is to prepare a presentation for next week and work on your project.

Schedule

May 6 Fifth presentation
May 13 Give a presentation and make final version available to me

(I will be gone May 5–6 (at IDA), but Grigor will lead class those two days.)